OpenStudy (anonymous):

Determine if the series converges or diverges: Sum[1/n^2-2n], n=3 to inf. My teacher told me it is telescoping, but im not seeing it.

OpenStudy (anonymous):

you wont see the telescope until you write is as partial fractions

OpenStudy (anonymous):

\[\frac{1}{n(n-2)}=\frac{1}{2(n-2)}-\frac{1}{2n}\]

OpenStudy (anonymous):

now it telescopes right?

OpenStudy (anonymous):

*Sigh* Yes, thanks

OpenStudy (anonymous):

yw